OpenStudy (anonymous):

The attic of a house has the shape of a triangular prism. The triangular end of the attic has a base that is 12 feet across and a height of 8 feet. The length of the attic is 30 feet from front to back.What is the volume of the attic?

OpenStudy (anonymous):

V=1/2 *h* (prism base area)

OpenStudy (anonymous):

prism base area = 8*30=240 V=1/2*8*240=960 fe^3